# Koby Hayashi

### v1

@@ 6 Conclusion
We have presented two novel algorithms for computing SymNMF in a randomized way. These methods are suitable for large problems as demonstrated by our findings on two data sets. Comparing these two methods we believe that the Algorithm LAI-SymNMF method is appropriate when a high quality LAI can be computed quickly and captures the underlying NMF signal. In other cases Algorithm LvS-SymNMF can be used. Our proposed methods achieve significant speeds ups of 5$\times$ or more and produce quality solutions on downstream clustering tasks. Additionally our techniques are applicable to standard NMF formulations as well. Finally we presented a number of theoretical results that justify why and when our proposed algorithms perform well in practice. These results include an analysis of leverage score sampling for approximately solving Nonnegative Least Squares problems, sample complexity analysis for the Hybrid Sampling scheme from [27], and some guarantees for Algorithm LAI-SymNMF.
Appendix A Additional Material for LAI-SymNMF
A.1 Additional Material and Reference for LAI-NMF Theory
Proposition 6 explicitly combines Proposition 1 and Lemma 9 to give a probabilistic error bound for LAI-NMF with a randomized low-rank input from the Algorithm RRF.
Proposition 6.
Given a matrix ${\bm{\mathbf{{X}}}}\in\mathbb{R}^{m\times n}$ compute a low-rank approximation ${\bm{\mathbf{{X}}}}\approx{\bm{\mathbf{{Q}}}}_{X}{\bm{\mathbf{{B}}}}_{X}$ where ${\bm{\mathbf{{Q}}}}_{X}\in\mathbb{R}^{m\times l}$ and ${\bm{\mathbf{{B}}}}_{X}\in\mathbb{R}^{l\times n}$ from the RRF. Then by Theorem 11, for any $0<\delta\ll 1$, we have that
$\mu_{RRF}=\sqrt{\Big{(}\sum_{j=k+1}^{n}\sigma_{j}^{2}\Big{)}+kC_{\delta}^{2}\sigma^{2}_{k+1}\Big{(}\frac{\sigma_{k+1}}{\sigma_{k}}\Big{)}^{4q}}\geq\|{\bm{\mathbf{{X}}}}-{\bm{\mathbf{{Q}}}}_{X}{\bm{\mathbf{{B}}}}_{X}\|_{F}$
holds with probability $1-\delta$. Define $\{{\bm{\mathbf{{W}}}}^{*},{\bm{\mathbf{{H}}}}^{*}\}=\operatorname*{arg\,min}_{{\bm{\mathbf{{W}}}},{\bm{\mathbf{{H}}}}\geq 0}\|{\bm{\mathbf{{Q}}}}_{X}{\bm{\mathbf{{B}}}}_{X}-{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}\|_{F}$ and the optimal approximation error $\upsilon^{*}=\min_{\{{\bm{\mathbf{{W}}}},{\bm{\mathbf{{H}}}}\}\geq 0}\|{\bm{\mathbf{{X}}}}-{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}\|_{F}$. Then with probability $1-\delta$
$\upsilon^{*}\leq\|{\bm{\mathbf{{X}}}}-{\bm{\mathbf{{W}}}}^{*}({\bm{\mathbf{{H}}}}^{*})^{\sf T}\|_{F}\leq 2\mu_{RRF}+\upsilon^{*}$ (36)
as in Proposition 1.
A.2 Comparison between LAI-NMF and Compressed-NMF
@@
The products ${\bm{\mathbf{{Q}}}}_{{\bm{\mathbf{{X}}}}}^{\sf T}{\bm{\mathbf{{X}}}}$ and ${\bm{\mathbf{{Q}}}}_{{\bm{\mathbf{{X}}}}^{\sf T}}^{\sf T}{\bm{\mathbf{{X}}}}^{\sf T}$ can be computed once but ${\bm{\mathbf{{Q}}}}_{{\bm{\mathbf{{X}}}}}^{\sf T}{\bm{\mathbf{{H}}}}$ and ${\bm{\mathbf{{Q}}}}_{{\bm{\mathbf{{X}}}}}^{\sf T}{\bm{\mathbf{{W}}}}$ need be recomputed each iteration. In this way a smaller problem is solved at each iteration resulting in computational savings. This method can be used for SymNMF by applying it to Equation 3. Additionally, in the symmetric case one need call the RRF only once.
We relate Compressed-NMF to LAI-NMF in the following way. Consider a QB-decomposition computed using the RRF where ${\bm{\mathbf{{Q}}}}_{{\bm{\mathbf{{X}}}}}=\textsc{RRF}({\bm{\mathbf{{X}}}},k,\rho,\omega)$ and ${\bm{\mathbf{{B}}}}_{{\bm{\mathbf{{X}}}}}={\bm{\mathbf{{Q}}}}_{{\bm{\mathbf{{X}}}}}^{\sf T}{\bm{\mathbf{{X}}}}$. Using the ${\bm{\mathbf{{Q}}}}{\bm{\mathbf{{B}}}}$ decomposition for LAI-NMF in Equation 20, the update to ${\bm{\mathbf{{H}}}}$ and its Quadratic Program (QP) form are
@@
Taking the QP formulation for the Compressed NMF update for ${\bm{\mathbf{{H}}}}$, $\min_{{\bm{\mathbf{{H}}}}\geq 0}\|{\bm{\mathbf{{Q}}}}_{{\bm{\mathbf{{X}}}}}^{\sf T}\big{(}{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}-{\bm{\mathbf{{X}}}}\big{)}\|_{F}^{2}$, yields
@@
The term on the right, ${\bm{\mathbf{{H}}}}{\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{Q}}}}_{{\bm{\mathbf{{X}}}}}{\bm{\mathbf{{B}}}}_{{\bm{\mathbf{{X}}}}}$, is the same for both problems. The difference between the two methods comes from the fact that the Gram matrix in LAI-NMF is ${\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{W}}}}$ but the Gram in Compressed-NMF contains the projection ${\bm{\mathbf{{Q}}}}_{{\bm{\mathbf{{X}}}}}{\bm{\mathbf{{Q}}}}_{{\bm{\mathbf{{X}}}}}^{\sf T}$. Empirically, for SymNMF, we find that these methods perform nearly identically. From this perspective we argue the affect of ${\bm{\mathbf{{Q}}}}_{{\bm{\mathbf{{X}}}}}{\bm{\mathbf{{Q}}}}_{{\bm{\mathbf{{X}}}}}^{\sf T}$ in the Gram matrix is minimal. Overall we propose that the LAI approach is easier to reason about, as it simply fits an NMF model of the low-rank approximation of ${\bm{\mathbf{{X}}}}$. As previously stated, using the basis ${\bm{\mathbf{{Q}}}}_{{\bm{\mathbf{{X}}}}}$ as a sketch for NLS problems is not justified theoretically.
A.3 Low-rank Approximate Input Projected Gauss-Newton with Conjugate Gradients for SymNMF
@@
Algorithm LAI-PGNCG-SymNMF : SymNMF of a Low-Rank Approximation
Appendix B Derivation of Symmetrically Regularized HALS
Now apply the standard HALS update
Finally the efficient updating method is derived as
The final update rule is written as
@@ C.1 Residual Checks
$\frac{\|{\bm{\mathbf{{X}}}}-{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}\|_{F}}{\|{\bm{\mathbf{{X}}}}\|_{F}}.$ (37)
@@
$\frac{\|{\bm{\mathbf{{U}}}}{\bm{\mathbf{{V}}}}^{\sf T}-{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}\|_{F}}{\|{\bm{\mathbf{{U}}}}{\bm{\mathbf{{V}}}}^{\sf T}\|_{F}}.$ (38)
@@
C.2 Projected Gradients
The projected gradient norm is defined as
$\sqrt{\|\nabla f_{W}\|_{F}^{2}+\|\nabla f_{H}\|_{F}^{2}}$ (39)
where the partial gradients are
$\nabla f_{W}=2({\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}{\bm{\mathbf{{H}}}}-{\bm{\mathbf{{A}}}}{\bm{\mathbf{{H}}}})$ (40)
$\nabla f_{H}=2({\bm{\mathbf{{H}}}}{\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{W}}}}-{\bm{\mathbf{{A}}}}^{\sf T}{\bm{\mathbf{{W}}}})$ (41)
and the projected gradient is
$(\nabla^{p}f_{W})_{ij}=\begin{cases}(\nabla f_{W})_{ij}\ \text{if}\ (\nabla f_{W})_{ij}<0\ \text{or}\ {\bm{\mathbf{{W}}}}_{ij}>0\\ 0\end{cases}$ (42)
similar for the partial gradient with respect to ${\bm{\mathbf{{H}}}}$. One issue with using the projected gradient as a comparison measure between algorithms is that different diagonal scalings of ${\bm{\mathbf{{W}}}}$ and ${\bm{\mathbf{{H}}}}$ can give different projected gradient values [23, 24]. However this is not an issue with the SymNMF objective. The gradient of Equation 2 is given by
$\nabla f_{H}=4({\bm{\mathbf{{H}}}}{\bm{\mathbf{{H}}}}^{\sf T}-{\bm{\mathbf{{X}}}}){\bm{\mathbf{{H}}}}$ (43)
Appendix D Additional Experimental Data
D.1 World of Science Data Set
This section contains some additional experiments on the World of Science data set. The first set of experiments are concerned with varying the $\rho$ or the column over sampling parameter in the RRF. The observation here is that increasing $\rho$ does not seem to have a beneficial impact on clustering quality or final residual. Figure 4 shows the runs for the various algorithms with $\rho=40$ and $\rho=80$. Section D.1 and Section D.1 give various metrics associated with these runs. Each algorithm was run 20 times. “Iters” is average number of iterations, “Time” is the average time in seconds, “Avg. Min-Res” is the average minimum residual achieved, “Min-Res” is the overall minimum residual achieved, and “Mean-ARI” is the mean of the ARI scores for each algorithm.
(a) BPP, p=40
(b) HALS, p=40
(c) GNCG, p=40
(d) BPP, p=80
(e) HALS, p=80
(f) GNCG, p=80
Figure 4: Normalized Residual error value for various SymNMF Algorithms on the WoS data set using a EDVW Hypergraph Representation.
[…]
The second set of experiments are run with $q=2$ and without the use of Algorithm Ada-RRF. Section D.1 gives data for this run. One can observe that method without IR do not achieve high quality residual or ARI scores. While IR can help fix these issues it does so at a higher computational cost. The convergence plots for these experiments are in Figures 4(c), 4(a) and 4(b).
(a) BPP
(b) HALS
(c) PGNCG
[…]
Table 6: Data Table for the WoS data set. Each algorithm was run 10 times. For LIA methods $q=2$, no Algorithm Ada-RRF is used. The columns record average number of iterations until the stopping criteria is met, average time in seconds, average minimum residual, and minimum residual achieved over all runs.
WoS Experiments System Details
D.2 Microsoft Open Academic Graph Experiments
This section contains additional data related to the Microsoft OAG experiments. Section D.2 shows the top 10 words for each cluster found by the HALS algorithm and Section D.2 shows the top 10 words for each cluster found by the LVS-HALS algorithm.
System Details
[…]
Appendix E Fast Residual Evaluation in NMF
@@
Appendix F Adaptive Randomized Range Finder
@@
Appendix G Update() Function
In this section we thoroughly explain the Update() function we use to simplify our pseudocode. The Framework for Alternating Updating NMF (FAUN) was proposed for the design of a massively parallel NMF code [21]. In the FAUN the matrix products ${\bm{\mathbf{{X}}}}^{\sf T}{\bm{\mathbf{{W}}}}$, ${\bm{\mathbf{{X}}}}{\bm{\mathbf{{H}}}}$, ${\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{W}}}}$, and ${\bm{\mathbf{{H}}}}^{\sf T}{\bm{\mathbf{{H}}}}$ are computed and used to perform updates. Many of the most effective NMF update rules require the computation of these 4 matrices as they appear in the gradient of the NMF NLS subproblems
$\min_{{\bm{\mathbf{{W}}}}\geq 0}\|{\bm{\mathbf{{X}}}}-{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}\|_{F}$ (44)
@@ Appendix G Update() Function
$\min_{{\bm{\mathbf{{H}}}}\geq 0}\|{\bm{\mathbf{{X}}}}-{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}\|_{F}.$ (45)
@@
Appendix H Theorems for Randomized Numerical Linear Algebra
@@
H.1 Structural Condition Theorems
We include statements of the two Structural Conditions taken from the work by Larsen and Kolda [27]. These theorems are included for completeness and reference.
Teorema 7.
@@ H.1 Structural Condition Theorems
$\|{\bm{\mathbf{{I}}}}-{\bm{\mathbf{{U}}}}^{\sf T}{\bm{\mathbf{{S}}}}^{\sf T}{\bm{\mathbf{{S}}}}{\bm{\mathbf{{U}}}}\|_{2}<\epsilon.$ (46)
H.2 Randomized Matrix Multiply
Teorema 8.

### v2

@@ 6 Conclusion
We have presented two novel algorithms for computing SymNMF in randomized ways. These methods are suitable for large problems as demonstrated by our findings on two data sets. Our proposed methods achieve significant speed ups of 5$\times$ or more and produce quality solutions on downstream clustering tasks. Additionally our techniques are applicable to standard NMF formulations as well. Finally we presented a number of theoretical results that justify why and when our proposed algorithms perform well in practice. These results include an analysis of leverage score sampling for approximately solving Nonnegative Least Squares problems, sample complexity analysis for the Hybrid Sampling scheme from [37], and some guarantees for Algorithm LAI-SymNMF.
Comparing these two methods we believe that the Algorithm LAI-SymNMF method is appropriate when a high quality LAI can be computed quickly and captures the underlying NMF signal. This is highlighted by Equation 17 which bounds the potential quality of the Algorithm LAI-SymNMF in terms of the quality of the Low-rank Approximate Input (LAI). In other cases Algorithm LvS-SymNMF can be used, particularly in the case of sparse input data, as it does not compute an upfront LAI but instead samples new subsets of rows at every iteration. Comparing these methods computationally, Algorithm LAI-SymNMF costs roughly $O(qm^{2}l)$ for computing the LAI and then $O(mkl)$ per iteration, where $l=O(k)$ is the rank of the intermediate approximation, and Algorithm LvS-SymNMF costs roughly $O(mks)=O(mk^{2}\,\max(\log(\frac{k}{\delta}),\frac{1}{\delta\epsilon_{r}}))$ per iteration, where $s$ is the number of row samples. The upfront cost of Algorithm LAI-SymNMF is expensive but its per iteration cost is lower than that of Algorithm LvS-SymNMF. This is assuming one stays in the regime of $k\ll m$. Additionally it is worth noting that both methods still require operations costing $O(mk^{2})$ for the products ${\bm{\mathbf{{H}}}}^{\sf T}{\bm{\mathbf{{H}}}}$ and ${\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{W}}}}$ or computing a thin-QR of ${\bm{\mathbf{{W}}}}$ and ${\bm{\mathbf{{H}}}}$. This again emphasizes the reliance of $k\ll m$ for computational efficiency. It is also important to note that the computational complexity costs of each algorithm do not tell the whole story. For example on large dense problems Algorithm LvS-SymNMF’s need to copy rows of the data matrix can be expensive as mentioned in Section 5.
Appendix A Derivation of Symmetrically Regularized HALS
Now apply the standard HALS update
Finally the efficient updating method is derived as
The final update rule is written as
Appendix B Additional Material for LAI-SymNMF
B.1 Comparison between LAI-NMF and Compressed-NMF
@@
The products ${\bm{\mathbf{{Q}}}}_{{\bm{\mathbf{{X}}}}}^{\sf T}{\bm{\mathbf{{X}}}}$ and ${\bm{\mathbf{{Q}}}}_{{\bm{\mathbf{{X}}}}^{\sf T}}^{\sf T}{\bm{\mathbf{{X}}}}^{\sf T}$ can be computed once but ${\bm{\mathbf{{Q}}}}_{{\bm{\mathbf{{X}}}}}^{\sf T}{\bm{\mathbf{{H}}}}$ and ${\bm{\mathbf{{Q}}}}_{{\bm{\mathbf{{X}}}}}^{\sf T}{\bm{\mathbf{{W}}}}$ need be recomputed each iteration. In this way a smaller problem is solved at each iteration resulting in computational savings. This method can be used for SymNMF by applying it to Eq. 3. Additionally, in the symmetric case one need call the RRF only once.
We relate Compressed-NMF to LAI-NMF in the following way. Consider a QB-decomposition computed using the RRF where ${\bm{\mathbf{{Q}}}}_{{\bm{\mathbf{{X}}}}}=\textsc{RRF}({\bm{\mathbf{{X}}}},k,\rho,\omega)$ and ${\bm{\mathbf{{B}}}}_{{\bm{\mathbf{{X}}}}}={\bm{\mathbf{{Q}}}}_{{\bm{\mathbf{{X}}}}}^{\sf T}{\bm{\mathbf{{X}}}}$. Using the ${\bm{\mathbf{{Q}}}}{\bm{\mathbf{{B}}}}$ decomposition for LAI-NMF in Eq. 14, the update to ${\bm{\mathbf{{H}}}}$ and its Quadratic Program (QP) form are
@@
Taking the QP formulation for the Compressed NMF update for ${\bm{\mathbf{{H}}}}$,
$\min_{{\bm{\mathbf{{H}}}}\geq 0}\|{\bm{\mathbf{{Q}}}}_{{\bm{\mathbf{{X}}}}}^{\sf T}\big{(}{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}-{\bm{\mathbf{{X}}}}\big{)}\|_{F}^{2}$, yields
@@
The term on the right, ${\bm{\mathbf{{H}}}}{\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{Q}}}}_{{\bm{\mathbf{{X}}}}}{\bm{\mathbf{{B}}}}_{{\bm{\mathbf{{X}}}}}$, is the same for both problems. The difference between the two methods comes from the fact that the Gram matrix in LAI-NMF is ${\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{W}}}}$ but the Gram in Compressed-NMF contains the projection ${\bm{\mathbf{{Q}}}}_{{\bm{\mathbf{{X}}}}}{\bm{\mathbf{{Q}}}}_{{\bm{\mathbf{{X}}}}}^{\sf T}$. Empirically, for SymNMF, we find that these methods perform nearly identically. From this perspective we argue the effect of ${\bm{\mathbf{{Q}}}}_{{\bm{\mathbf{{X}}}}}{\bm{\mathbf{{Q}}}}_{{\bm{\mathbf{{X}}}}}^{\sf T}$ in the Gram matrix is minimal. Overall we propose that the LAI approach is easier to reason about, as it simply fits an NMF model of the low-rank approximation of ${\bm{\mathbf{{X}}}}$. As previously stated, using the basis ${\bm{\mathbf{{Q}}}}_{{\bm{\mathbf{{X}}}}}$ as a sketch for NLS problems is not justified theoretically.
B.2 Low-rank Approximate Input Projected Gauss-Newton with Conjugate Gradients for SymNMF
@@ C.1 Residual Checks
(32) $\frac{\|{\bm{\mathbf{{X}}}}-{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}\|_{F}}{\|{\bm{\mathbf{{X}}}}\|_{F}}.$
@@
(33) $\frac{\|{\bm{\mathbf{{U}}}}{\bm{\mathbf{{V}}}}^{\sf T}-{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}\|_{F}}{\|{\bm{\mathbf{{U}}}}{\bm{\mathbf{{V}}}}^{\sf T}\|_{F}}.$
@@
C.2 Fast Residual Evaluation in NMF
@@
C.3 Projected Gradients
The projected gradient norm is defined as
(34) $\sqrt{\|\nabla f_{W}\|_{F}^{2}+\|\nabla f_{H}\|_{F}^{2}}$
where the partial gradients are
(35) $\nabla f_{W}=2({\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}{\bm{\mathbf{{H}}}}-{\bm{\mathbf{{A}}}}{\bm{\mathbf{{H}}}})$
(36) $\nabla f_{H}=2({\bm{\mathbf{{H}}}}{\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{W}}}}-{\bm{\mathbf{{A}}}}^{\sf T}{\bm{\mathbf{{W}}}})$
and the projected gradient is
(37) $(\nabla^{p}f_{W})_{ij}=\begin{cases}(\nabla f_{W})_{ij}\ \text{if}\ (\nabla f_{W})_{ij}<0\ \text{or}\ {\bm{\mathbf{{W}}}}_{ij}>0\\ 0\end{cases}$
similar for the partial gradient with respect to ${\bm{\mathbf{{H}}}}$. One issue with using the projected gradient as a comparison measure between algorithms is that different diagonal scalings of ${\bm{\mathbf{{W}}}}$ and ${\bm{\mathbf{{H}}}}$ can give different projected gradient values [32, 33]. However this is not an issue with the SymNMF objective. The gradient of Eq. 2 is given by
(38) $\nabla f_{H}=4({\bm{\mathbf{{H}}}}{\bm{\mathbf{{H}}}}^{\sf T}-{\bm{\mathbf{{X}}}}){\bm{\mathbf{{H}}}}$
Appendix D Adaptive Randomized Range Finder
@@
Appendix E Update() Function
In this section we thoroughly explain the Update(${\bm{\mathbf{{G}}}},{\bm{\mathbf{{Y}}}}$) function we use to simplify our pseudocode. This function takes in two matrices ${\bm{\mathbf{{G}}}}$ which is $k\times k$ and ${\bm{\mathbf{{Y}}}}$ which is $k\times m$ or $k\times n$. The Framework for Alternating Updating NMF (FAUN) was proposed for the design of a massively parallel NMF code [29]. In the FAUN the matrix products ${\bm{\mathbf{{X}}}}^{\sf T}{\bm{\mathbf{{W}}}}$, ${\bm{\mathbf{{X}}}}{\bm{\mathbf{{H}}}}$, ${\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{W}}}}$, and ${\bm{\mathbf{{H}}}}^{\sf T}{\bm{\mathbf{{H}}}}$ are computed and used to perform updates. Many of the most effective NMF update rules require the computation of these 4 matrices as they appear in the gradient of the NMF NLS subproblems
(39) $\min_{{\bm{\mathbf{{W}}}}\geq 0}\|{\bm{\mathbf{{X}}}}-{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}\|_{F}$
@@ Appendix E Update() Function
(40) $\min_{{\bm{\mathbf{{H}}}}\geq 0}\|{\bm{\mathbf{{X}}}}-{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}\|_{F}.$
@@
Appendix F Theorems for Randomized Numerical Linear Algebra
@@
F.1 Structural Condition Theorems
We include statements of the two Structural Conditions taken from the work by Larsen and Kolda [37]. Prior versions and similar statements of this result are originally from [56, 41] These theorems are included for completeness and reference.
Theorem F.1.
@@ F.1 Structural Condition Theorems
(41) $\|{\bm{\mathbf{{I}}}}-{\bm{\mathbf{{U}}}}^{\sf T}{\bm{\mathbf{{S}}}}^{\sf T}{\bm{\mathbf{{S}}}}{\bm{\mathbf{{U}}}}\|_{2}<\epsilon.$
F.2 Randomized Matrix Multiply
For reference we state a result originally from [12].
Theorem F.2.